OpenStudy (anonymous):

I know that 0=(20.0m/s)t - 4.90m/s^2)t^2 =t(20.0m/s - 4.90m/s^t) is equal to t = 4.08 but how do you do that?

OpenStudy (anonymous):

0 = 20t - 4.9t factor 0 = t(20 - 4.9t) split into 2 equations (b/c zero product property) t = 0 20 - 4.9t = 0 add 4.9t to both sides 20 = 4.9t divide by 4.9 4.08 = t

OpenStudy (anonymous):

ohh ok thanks